# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@ 4.3 Softmax Activation
Building on Definition 4.1, we develop a novel algorithm to answer differentially private weighted Softmax queries using the polynomial kernel method from AS [6]. Specifically, in Eq.(2), there is a term that computes the weighted $\ell_{2}^{2}$ distance, which we calculate using DPTreeHighDim. We then compute the exact term for the weighted $\ell_{2}^{2}$ norms of the approximation kernel. By summing these terms with a controlled error, we extend DPTreeHighDim to answer the Softmax query efficiently. More details can be found in Section I.
Theorem 4.2 (Softmax query, informal version of Theorem I.8).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Our data structure DPTreeSoftmax (Algorithm 3) uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1))$. (Algorithm 3) It takes $O(nr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
@@
    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      $+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
Remark 4.3.
In Theorem 4.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section C.5. Besides, note that the error bound in Theorem 4.2 does not depend on $\delta$ but depends on $\delta^{\prime}$. The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ reduces the error by a small constant (Fact 2.5). In practice, we set $\delta$ as a small positive constant close to 00. Please refer to the Lemma D.6 for more details.
@@ 4.4 Adaptive Query Data Structure
We adapt our DPTreeSoftmax to DPTreeSoftmaxAdaptive (Algorithm 1) to solve the adaptive query problem. By proving it can handle any query within the query space with a certain error, we ensure it effectively processes adaptive queries. We first boost the constant probability to high probability using the Chernoff bound (Lemma B.2). Employing an $\epsilon_{0}$-net argument and the union bound, we bound all query points within the net. Finally, we use the Lipschitz property of the weighted Softmax distance function with an additive error to bound all points in the query space. The corresponding proofs can be found in Section H and Section I.
Theorem 4.4 (Adaptive query Softmax data structure, informal version of Theorem I.11).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnr)$ spaces to solve the weighted Softmax query problem for the dataset $X\subset[0,R]^{d}$ and supports the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$. It takes $O(lnr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output a number $z$ such that
@@
    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      $+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
Remark 4.5.
We describe the parallelization of our algorithms. In the second for loop of Init and the for loop of DistanceQuery in Algorithm 3, the $r$ DPTreeDistance data structures instantiated for each coordinate are independent of each other. In addition, the for loops in Algorithm 1 are also parallelizable since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent. After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\alpha^{-1}\log^{2}n)$ in Algorithm 1 with $O(lr)$ GPU process.
@@ 5 Discussion
How do we extend to self-attention?
As self-attention is a more fundamental module in LGMs, we would like to extend our data structure to this setting. However, the challenge we faced was the dynamic update in tree nodes for each query for self-attention, which our current analysis does not support. How we can solve this challenge is crucial, and we leave it as our future direction.
Why not add noise to some other places?
Where and how to add DP noises is an important problem to ask during the DP algorithm design. In this paper, we consider the problem of $\sum_{i=1}^{n}w_{i}\exp(\langle x_{i},y\rangle/d)$ where $y,x_{i}\in[0,R]^{d}$ and $w\in[-R_{w},R_{w}]^{n}$ (Definition 4.1). Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorihtm 2). From Lemma C.3 and the way we initialize DPTree in Algorithm 5, we see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
A simple method for adding noise involves adding $n$ noises to a length $n$ array, with each item $w_{i}\exp(\langle x_{i},y\rangle/d)$ for $i\in[n]$. However, this approach increases the error by a factor of $n$ by basic composition (Fact B.6) and also makes the model dependent on the number of queries. Besides, it only supports a single query and requires rebuilding the tree for each new query, rendering it impractical. In contrast, our current noise-adding technique (Lines 9 and 15 of Algorithm 2) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$. This method also supports multiple queries, eliminating the need to rebuild the tree each time.
How to remove the relative error parameter $\alpha$?
The relative error parameter $\alpha$ in Theorem 3.1 appears because of the $(1+\alpha)$-approximation introduced in Algorithm 5 (Remark F.3) to reduce the number of required iterations from naive $O(n)$ to $O(\log(n)/\alpha)$. However, we notice that a recent work [59] does not utilize $(1+\alpha)$-approximation and still achieves $O(\log n)$ iteration number. They introduce a new tree node representation where each node stores the sum of distances from one point to multiple points, enabling the answer to be divided into only $\log n$ values, each combining two distance values, two count values, and $y$ itself. Our DPTree algorithms can be integrated with their method, thus removing parameter $\alpha$.
@@ 6 Conclusion
To our knowledge, we are the first work to provide differential privacy for cross-attention. This paper presents the DPTree data structures, which provide a differential privacy guarantee for the cross-attention module in large generative models. This is achieved by transforming the cross-attention mechanism into a weighted distance problem. Furthermore, our algorithm is robust to adaptive queries, allowing users to interact with the model arbitrarily without extracting sensitive information from the system prompts or RAG data. Our results may inspire more privacy algorithm design in large generative models.
Acknowledgement
Research is partially supported by the National Science Foundation (NSF) Grants 2023239-DMS, CCF-2046710, and Air Force Grant FA9550-18-1-0166.
Appendix
@@ Acknowledgement
The appendix is organized as follows. In Section A, we provide more related works. In Section B, we give the preliminary of our paper. In Section C, we offer an outline of our proof techniques. In Section D, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee. In Section E, we show how to solve weighted distance problem. In Section F, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{1}$ distance problem with DP and accuracy guarantee. In Section G, we present the analysis of our DPTreeHighDim (Algorithm 7) data structure, which can address the high-dimensional $\ell_{1}$ distance problem while ensuring differential privacy and accuracy guarantees. In Section H, we show how we can handle adaptive query. In Section I, we show how to extend our algorithm to Softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 1).
Appendix A More Related Work
Differential Privacy Guarantee Analysis.
Ever since DMNS [19] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [22, 67, 115, 84, 111]. DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [50, 24, 3, 61, 51, 37, 10, 12, 25, 16, 46, 78, 79, 56, 62, 28, 29, 60, 17, 13, 18, 27, 40, 39, 64, 38, 23, 95, 21, 107, 94, 42, 70, 66]. Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [20]. For example, truncated Laplace mechanism [32] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
Cross-Attention in System Prompt, RAG, Stable Diffusion and More.
Cross-attention [97], first introduced in language translation, is a widely used technique in many advanced AI systems. For example, Stable Diffusion [87, 73, 98, 103, 106] and SORA [81] employ cross-attention as a core module for a text-to-image conditional generation. This technique is also utilized by other multimodal models [72], including Imagen [90] and Diffusion Transformer [85]. In the realm of text-to-image editing, HMT+ [49] analyzes and controls the cross-attention module to enable editing without requiring additional training. Furthermore, YYB+ [114] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas. In addition, Retrieval Augmented Generation (RAG) [68, 11, 43], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module. Cross-attention also has other applications. ORST [82] demonstrates that the prompt-tuning [71] task can be formulated as cross-attention, while CFP [14] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation. Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [104], such as spatial reasoning [101], mathematical reasoning [65], in-context learning ability [93], compositional ability [109], few-shot adaptation ability [89, 110], and so on. There are some other works that used cross attention in Hopfield Models [52, 100, 48, 108, 99, 44, 45].
Appendix B More Preliminary
In Section B.1, we give the probability tools we use in the paper. In Section B.2, we provide the algebraic facts we use. In Section B.3, we give the DP facts we use in the paper. In Section B.4, we compare between popular DP mechanisms.
B.1 Probability Tools
@@
Lemma B.1 (Markov’s inequality).
@@
Lemma B.2 (Chernoff bound, [15]).
Let $x_{i}$ be a Bernoulli random variable with probability $p_{i}$ of being equal to $1$ and $1-p_{i}$ of being equal to 00, and all $x_{i}$ for $i\in[n]$ are independent. Let $x=\sum_{i=1}^{n}x_{i}$. Let $\mu=\operatorname*{{\mathbb{E}}}[x]=\sum_{i=1}^{n}p_{i}$. Then, for all $\delta>0$ we have
@@
Lemma B.3 (Chebyshev’s inequality).
@@
B.2 Algebraic Facts
Fact B.4 (Upper bound of exponential, Fact C.9 in LSSZ24b [73]).
@@ B.2 Algebraic Facts
B.3 DP Facts
In this section, we present several facts about differential privacy (DP). We first state the post-processing property, which means, in an algorithm, if one step is DP, all the following steps are DP.
Fact B.5 (Post-processing, see Fact 2.1 in GKK+ [37]).
@@ B.3 DP Facts
Fact B.6 (Basic composition, see Fact 2.3 in GKK+ [37]).
@@
Fact B.7 (Parallel composition, see Fact 2.4 in GKK+ [37]).
@@
Theorem B.8 (Advanced composition, see Theorem 3.20 in DR [20]).
@@
B.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms

### v3

@@ 5.3 Softmax Activation
Building on Definition 5.1, we develop a novel algorithm to answer differentially private weighted Softmax queries using the polynomial kernel method from [4]. Specifically, in Eq.(5.1), the three terms compute the weighted $\ell_{2}^{2}$ distance, which we calculate using DPTreeHighDim. By summing these terms with a controlled error, we extend DPTreeHighDim to answer the Softmax query efficiently. More details can be found in Section G.
Theorem 5.2 (Softmax query, informal version of Theorem G.7).
Let $R\geq 1$. Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$. Let $r={2s+2d\choose 2s}$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Our data structure DPTreeSoftmax (Algorithm 3) uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
[…]
@@
    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
In Theorem 5.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section G. Besides, note that the error bound in Theorem 5.2 does not depend on $\delta$ but depends on $\delta^{\prime}$. The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ reduces the error by a small constant (Fact 3.5). In practice, we set $\delta$ as a small positive constant close to 0. Please refer to the Lemma B.7 for more details.
@@ 5.4 Adaptive Query Data Structure
We adapt our DPTreeSoftmax to DPTreeSoftmaxAdaptive (Algorithm 7) to solve the adaptive query problem. By proving it can handle any query within the query space with a certain error, we ensure it effectively processes adaptive queries. We first boost the constant probability to high probability using the Chernoff bound (Lemma A.2). Employing an $\epsilon_{0}$-net argument and the union bound, we bound all query points within the net. Finally, we use the Lipschitz property of the weighted Softmax distance function with an additive error to bound all points in the query space. The corresponding proofs can be found in Section F and Section G.
Theorem 5.3 (Adaptive query Softmax data structure, informal version of Theorem G.10).
Let $R\geq 1$. Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $r={2s+2d\choose 2s}$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f}))).$.
There is a data structure DPTreeSoftmaxAdaptive (Algorithm 7) that uses $O(lnr)$ spaces to solve the weighted Softmax query problem for the dataset $X\subset[0,R]^{d}$ and supports the following operations:
[…]
@@
    • we have $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
We describe the parallelization of our algorithms. In the second for loop of Init and the for loop of DistanceQuery in Algorithm 3, the $r$ DPTreeDistance data structures instantiated for each coordinate are independent of each other. In addition, the for loops in Algorithm 7 are also parallelizable since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent. After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\log n)$ in Algorithm 7 with $O(lr)$ GPU process.
@@ 6 Discussion
How do we extend to self-attention and other data structures? As self-attention is a more fundamental module in LGMs, we would like to extend our data structure to this setting. However, the challenge we faced was the dynamic update in tree nodes for each query for self-attention, which our current analysis does not support. How we can solve this challenge is crucial, and we leave it as our future direction. Moreover, we observe that [52] introduces the DP matrix mechanism, which offers an alternative to our currently used binary tree data structure. A preliminary idea for extending this is as follows: consider $A=\exp(QK^{\top}/d)$ as defined in Definition 1.1, where $Q$ of size $m\times d$ represents the query matrix with $m$ linear queries, and $K$ serves as the database. Leveraging the results from [52], we could design an alternative algorithm to enhance the current binary tree data structure, DPTree. We leave this exploration for future work.
Why not add noise to some other places? Where and how to add DP noises is an important problem to ask during the DP algorithm design. In this paper, we consider the problem of $\sum_{i=1}^{n}w_{i}\exp(\langle x_{i},y\rangle/d)$ where $y,x_{i}\in[0,R]^{d}$ and $w\in[-R_{w},R_{w}]^{n}$ (Definition 5.1). Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorithm 2). From Lemma B.3 and the way we initialize DPTree in Algorithm 4, we see that the sensitivity $\Delta$ of this problem is $2R_{w}$. A simple method for adding noise involves adding $n$ noises to a length $n$ array, with each item $w_{i}\exp(\langle x_{i},y\rangle/d)$ for $i\in[n]$. However, this approach increases the error by a factor of $n$ by basic composition (Fact A.8) and also makes the model dependent on the number of queries. Besides, it only supports a single query and requires rebuilding the tree for each new query, rendering it impractical. In contrast, our current noise-adding technique (Lines 8 and 14 of Algorithm 2) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$. This method also supports multiple queries, eliminating the need to rebuild the tree each time.
@@ 7 Conclusion
To the best of our knowledge, this paper represents the first attempt to incorporate differential privacy into the cross-attention mechanisms. This paper presents the DPTree data structures, which provide a differential privacy guarantee for the cross-attention module in large generative models. This is achieved by transforming the cross-attention mechanism into a weighted distance problem. Furthermore, our algorithm is robust to adaptive queries, allowing users to interact with the model arbitrarily without extracting sensitive information from the system prompts or RAG data. Our results may inspire more privacy algorithm design in large generative models.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Not Applicable]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Not Applicable]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Not Applicable]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Not Applicable]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@
The appendix is organized as follows. In Section A, we give the preliminary of our paper. In Section B, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee. In Section C, we show how to solve weighted distance problem. In Section D, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{p}^{p}$ distance problem with DP and accuracy guarantee. In Section E, we present the analysis of our DPTreeHighDim (Algorithm 6) data structure, which can address the high-dimensional $\ell_{p}^{p}$ distance problem while ensuring differential privacy and accuracy guarantees. In Section F, we show how we can handle adaptive query. In Section G, we show how to extend our algorithm to Softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 7).
Appendix A More Preliminary
In Section A.1, we give the probability tools we use in the paper. In Section A.2, we provide the algebraic facts we use. In Section A.3, we give the DP facts we use in the paper. In Section A.4, we compare between popular DP mechanisms.
A.1 Probability Tools
@@
Lemma A.1 (Markov’s inequality).
@@
Lemma A.2 (Chernoff bound, [13]).
Let $x_{i}$ be a Bernoulli random variable with probability $p_{i}$ of being equal to $1$ and $1-p_{i}$ of being equal to 0, and all $x_{i}$ for $i\in[n]$ are independent. Let $x=\sum_{i=1}^{n}x_{i}$. Let $\mu=\operatorname*{{\mathbb{E}}}[x]=\sum_{i=1}^{n}p_{i}$. Then, for all $\delta>0$ we have
@@
Lemma A.3 (Chebyshev’s inequality).
@@
A.2 Algebraic Facts
Fact A.4 (Upper bound of exponential, Fact C.9 in [57]).
@@ A.2 Algebraic Facts
A.3 DP Facts
In this section, we present several facts about differential privacy (DP).
We first define vector neighboring dataset and sensitivity.
Definition A.5 (Vector neighboring dataset).
Definition A.6 (Vector sensitivity).
We state the post-processing property, which means, in an algorithm, if one step is DP, all the following steps are DP.
Fact A.7 (Post-processing, see Fact 2.1 in [27]).
@@ A.3 DP Facts
Fact A.8 (Basic composition, see Fact 2.3 in [27]).
@@
Fact A.9 (Parallel composition, see Fact 2.4 in [27]).
@@
Theorem A.10 (Advanced composition, see Theorem 3.20 in [17]).
@@
A.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms